# Meisam Kabiri1, and Holger Voos12 1 Interdisciplinary Center for Security Reliability and Trust (SnT)

### v3

@@ 78 GHz ToA Failed Failed
Figure 9: RGB-D trajectory and map point estimates in the global frame for Dataset0 with 78 GHz ToA integration. The SLAM results show keyframe pose estimates (blue lines), mapped features (black dots), and the ground truth trajectory (red line). Coordinate axes are color-coded: red for X, green for Y, and blue for Z. The trajectory estimate, enhanced with 78 GHz ToA data, aligns closely with the ground truth in the global frame, demonstrating the system’s ability to maintain accurate global positioning without post-processing alignment.
V-B Analysis of SLAM Modes with ToA Integration
V-B1 RGB-D and RGB-D Inertial Modes
For RGB-D SLAM mode, the baseline achieves exceptional local accuracy. Integrating 78 GHz ToA measurements shows slight improvements in most cases, while 28 GHz ToA integration results in minor degradation. This performance difference stems from two factors:
  • Baseline Superiority and 6-DoF Interpretation: the baseline’s local accuracy already exceeds ToA measurement precision as shown in Table I. Additionally, ToA data offers only distance information, and translating these measurements into full 6-DoF pose estimations can inherently limit accuracy.
  • Post-Hoc Alignment and Fusion Complexity: Post-hoc alignment compensates for global positioning errors, minimizing the impact of ToA integration on local accuracy. Furthermore, fusing additional sensor data introduces complexities and potential calibration challenges, which can lead to minor inconsistencies in the local ATEs.
For global evaluation, ToA integration effectively enables global localization, with 78 GHz measurements achieving global ATEs of approximately 0.1m on Dataset0, showcasing excellent performance and demonstrating robust global localization capabilities.
In the RGB-D Inertial mode, a similar pattern emerges. ToA integration offers modest improvements in local accuracy while enabling efficient global SLAM, addressing the limitations of the baseline, which lacks external references for global positioning.
V-B2 Monocular (MONO) Modes
For MONO SLAM mode, the baseline encounters fundamental challenges due to scale ambiguity and the absence of depth information, leading to high ATEs and complete failure on Dataset 0. Even with post-hoc alignment, local errors remain above 1m, highlighting the limitations of the baseline approach.
The integration of ToA offers critical scale and positioning information, resulting in substantial performance improvements. At 78 GHz, global ATEs achieve an accuracy of 0.121m on Dataset 1. Local performance also benefits significantly, primarily due to enhanced scale and transformation estimation during SLAM operations, enabling more robust and reliable localization.
V-B3 Monocular Inertial Modes
For MONO Inertial mode, the baseline fails across all datasets due to the compounding effects of inertial sensor noise. ToA integration significantly enhances both accuracy and robustness by enabling reliable scale and global transformation estimation while stabilizing the fusion of inertial data. At 78 GHz, global ATE achieves 0.123m on Dataset 3, demonstrating the potential of ToA integration in improving performance.
However, challenges persist in certain scenarios. For example, Dataset 2 exhibits a high global ATE of 0.649m even with 78 GHz ToA data. This highlights that while ToA integration addresses many limitations, effectively managing multiple sensor modalities under diverse conditions remains a complex task requiring further optimization.
V-C Impact of 5G Network Frequency on ToA Performance
While the 78 GHz 5G network configuration generally enables better ToA-based localization performance compared to 28 GHz in the global evaluation, thanks to its greater bandwidth and subcarrier spacing, this pattern isn’t universal across datasets. Performance variations between the two configurations may arise from measurement noise and ToA estimation accuracy under different conditions.
In the local evaluation, differences between the two 5G network configurations are less pronounced due to the effects of post-hoc alignment. However, the improved precision of the 78 GHz configuration often results in better trajectory consistency in most scenarios.
V-D Real-Time Performance
Our approach achieves real-time operation on ROS bag files at normal playback speed, running on a laptop with Intel i9-10885H CPU (2.40GHz, 16 cores) and 32GB RAM under Ubuntu 20.04. This computational efficiency makes the system practical for real-time drone navigation applications.
V-E Summary and Key Findings
The integration of ToA measurements with various SLAM configurations reveals distinct effects on both global and local performance. Most importantly, ToA integration addresses a fundamental limitation of traditional SLAM systems by enabling global localization capabilities, which are otherwise unattainable without external references.
In minimal sensor configurations, such as Monocular SLAM, ToA integration proves transformative. Baseline Monocular SLAM suffers from scale ambiguity and significant trajectory errors, leading to local errors exceeding 1 meter even after scale correction. With ToA integration, the system achieves substantial improvements, with average local accuracies across datasets reaching 9.6 cm with 28 GHz and 6.3 cm with 78 GHz, notably outperforming the RGB-D SLAM baseline of 11.5 cm. Similarly, Monocular Inertial SLAM, which fails entirely in its baseline configuration, becomes fully operational with ToA integration, achieving average local accuracies around 11 cm across both frequency bands.
Conversely, for sensor-rich configurations like RGB-D SLAM, where the baseline already achieves high precision (around 11.5cm local error), ToA integration offers limited benefits for local accuracy. The 78 GHz configuration shows a modest 1.3% improvement, while 28 GHz integration slightly degrades performance by 5.2%. This pattern suggests that ToA measurements, while valuable for global positioning, may introduce additional noise that impacts local accuracy when combined with already precise depth measurements. An interesting observation is the superior performance of 78 GHz ToA integration compared to 28 GHz across most configurations, particularly in global accuracy. This aligns with the theoretical advantages of higher frequency bands offering greater bandwidth and improved measurement precision. However, the performance gap between frequencies becomes less pronounced in local accuracy evaluation due to the effects of post-hoc alignment.
These findings highlight that ToA integration’s impact is most beneficial in scenarios with minimal sensor configurations or when global positioning capabilities are required, while its utility for local accuracy may be limited in setups with high-precision depth sensing.
VI System Limitations and Future Enhancements
While the experimental results demonstrate the efficacy of integrating simulated 5G ToA measurements with traditional visual and inertial data for SLAM, several limitations must be acknowledged:
  • Ideal Line-of-Sight (LoS) Conditions in ToA Simulation: The current ToA simulation assumes ideal LoS conditions, which may not hold true in real-world scenarios. Non-Line-of-Sight (NLoS) conditions, caused by obstacles, can degrade ToA measurement accuracy and impact SLAM performance. Future work needs more realistic channel models for complex environments.
  • Fixed Measurement Uncertainties: The current approach uses static information matrices for ToA and visual-inertial measurements. Environmental factors can affect sensor reliability, requiring adaptive techniques that adjust measurement uncertainties dynamically.
  • Assumption of Fixed and Known Base Station Positions: The experiments assume that base station positions are fixed and precisely known. While this simplifies the implementation, practical deployments may involve scenarios where some base stations are mobile, have uncertain positions, or are entirely unknown. Importantly, if none of the base station positions are known, the system would lack a global reference frame, making global SLAM infeasible. However, in cases where some base station positions are known while others are dynamic or unknown, partial global localization might still be achievable. Future research should investigate robust techniques for joint localization and dynamic base station position estimation to address these challenges and extend the system’s applicability to more realistic scenarios.
  • Simulation-Based ToA Data Limitations: Simulated ToA data may not fully capture real-world signal propagation complexities, hardware noise, and environmental dynamics. Validation with experimental ToA measurements is crucial for assessing practical performance.
VII Conclusion
This paper presented a novel approach for integrating 5G ToA measurements with ORB-SLAM3 to enable globally consistent localization and mapping. Our method seamlessly embedded ToA data into the optimization framework, allowing real-time estimation of the transformation between local and global reference frames. Through extensive evaluation on five real-world datasets, we demonstrated that ToA integration maintained the high local accuracy of RGB-D SLAM while enabling robust global positioning. The approach proved particularly effective for monocular SLAM, resolving scale ambiguity and maintaining tracking in scenarios where standard implementations failed.
Future work should focus on addressing current limitations, including Non-Line-of-Sight conditions, adaptive measurement uncertainties, and scenarios involving partially known or dynamic base station positions. While the method relies on the assumption that some base stations have fixed and known positions to provide a global reference frame, exploring techniques for joint localization and dynamic base station position estimation could significantly expand its applicability. Additionally, incorporating real-world ToA measurements and refining the system for practical deployment challenges remain key areas of exploration.
The successful integration of 5G ToA with visual-inertial SLAM has opened new possibilities for robust indoor navigation, particularly in applications requiring precise global positioning, such as warehouse automation and drone navigation. By leveraging partial knowledge of base station positions and robust optimization frameworks, this approach represents a significant step toward scalable, globally consistent SLAM solutions.

### v5

@@ 78 GHz ToA Failed Failed
Figure 9: RGB-D trajectory and map point estimates in the global frame for AL01 with 78 GHz ToA integration. The SLAM results show keyframe pose estimates (blue lines), mapped features (black dots), and the ground truth trajectory (red line). Coordinate axes are color-coded: red for X, green for Y, and blue for Z. The trajectory estimate, enhanced with 78 GHz ToA data, aligns closely with the ground truth in the global frame, demonstrating the system’s ability to maintain accurate global positioning without post-processing alignment.
V-A1 Analysis of SLAM Modes with ToA Integration
The integration of ToA measurements into various SLAM modes reveals several key insights across both the Aerolab and EuRoC datasets. For RGB-D SLAM, the baseline already achieves good local accuracy (0.115 m average), while ToA integration shows inconsistent effects—78 GHz ToA yields a modest improvement of 1.3% in local accuracy, whereas 28 GHz ToA causes a 5.2% degradation. AL01 exhibits the most significant improvement (14.1%) with 78 GHz ToA, contrasting with AL05’s 12.3% degradation using the same configuration.
RGB-D Inertial SLAM follows a similar pattern with more pronounced variations: 78 GHz ToA shows an average local improvement of 5.97%, while 28 GHz ToA offers minimal improvement (0.42%). AL03 presents anomalous degradation with both ToA frequencies (-6.9% and -69.4%, excluded from averages), indicating potential trajectory-specific challenges.
In Monocular SLAM, where scale estimation is inherently problematic, 78 GHz ToA integration provides a substantial 11.4% improvement in local accuracy while maintaining excellent scale estimation (0.156% average error). Meanwhile, 28 GHz ToA integration results in a 7.8% degradation despite reasonable scale estimation (0.300% error). On the EuRoC dataset, both Stereo (8.2% improvement with 78 GHz) and Monocular modes (3.1% improvement with 78 GHz) demonstrate consistent benefits from 78 GHz ToA integration.
The most significant advantage of ToA integration lies in enabling global positioning across all SLAM modes—a capability absent in baseline configurations. The 78 GHz ToA consistently outperforms 28 GHz ToA in global accuracy: RGB-D (0.133 m vs. 0.198 m), RGB-D Inertial (0.262 m vs. 0.280 m), and Monocular (0.126 m vs. 0.267 m) on the Aerolab dataset. This performance disparity stems from the higher precision of 78 GHz measurements.
Environmental factors and trajectory characteristics significantly influence ToA integration performance, as evidenced by the inconsistent results across datasets. Dataset 2 particularly demonstrates how specific conditions can challenge ToA integration across multiple SLAM modes. Despite these variations, the addition of global positioning capabilities represents a substantial enhancement for applications requiring absolute positioning, even when local accuracy improvements remain marginal or inconsistent.
V-B SLAM Performance with Unknown Base Station Configurations
To evaluate SLAM performance with unknown base station locations, we modified our system architecture by changing the base station nodes from fixed to non-fixed status while maintaining a fixed transformation node. This configuration eliminates the requirement for prior knowledge of base station locations, though it does preclude global positioning capabilities. We further relaxed operational conditions by:
  • Reducing the number of active base stations
  • Implementing sequential line-of-sight (LOS) connections where only one base station maintains LOS during specific, non-overlapping time intervals
  • Limiting base station coverage to partial trajectory segments rather than the complete path
We evaluated these scenarios across both minimal and rich sensor configurations under 78 GHz 5G network. For minimal sensor setups, we focused on monocular configurations using both the Aerolab and EuRoC MAV datasets, as ToA integration typically shows the most pronounced effects with limited sensor data. For rich sensor configurations (RGB-D and stereo), we investigated ToA’s potential as an alternative to loop closure. Loop closure may be unavailable in many practical scenarios, such as:
  • Linear trajectories without path intersections
  • Exploration of previously unvisited areas
  • Time-critical operations that prevent revisiting locations
  • Environments with significant appearance changes
  • Resource-constrained systems where loop closure detection is computationally prohibitive
We identified sequences where loop closure was naturally occurring and contributing to the system’s accuracy (AL01, V102, V103, V202, and V203). To evaluate ToA measurements as a potential alternative to loop closure for drift correction, we deliberately disabled loop closure in these specific sequences. This approach allowed us to directly assess whether ToA integration could effectively compensate for the absence of loop closure in practical scenarios where visual loop closure might be unavailable or unreliable.
For comparison with recent literature, we also replicated the experimental setup from [41] using the EuRoC Machine Hall dataset. While their work used simulated ToA measurements from one base station at 30 Hz ($\sigma=5$ cm Gaussian noise, base station at [10,10,10]), we modified the scenario to operate at a more practical 5 Hz measurement frequency. This allows direct comparison while reflecting more realistic operational constraints.
V-B1 Sequential Unknown Base Station Operation (Monocular)
TABLE IV: MONO SLAM with Sequential Unknown Base Stations
Sequence Configuration Local (m) Local (Unscaled) (m) Scale Local Imp. V101 Baseline 0.0950 N/A N/A Sequential BS 0.0944 0.1222 3.81% 0.63% V102 Baseline 0.1004 N/A N/A Sequential BS 0.1001 0.1020 0.71% 0.30% V103 Baseline 0.1160 N/A N/A Sequential BS 0.1068 0.1144 1.63% 7.93% V201 Baseline 0.0860 N/A N/A Sequential BS 0.0848 0.0924 1.33% 1.39% V202 Baseline 0.1760 N/A N/A Sequential BS 0.1748 0.1803 0.38% 0.68% AL02 Baseline 0.0550 N/A N/A Sequential BS 0.0534 0.0709 0.90% 2.90% AL04 Baseline 0.0740 N/A N/A Sequential BS 0.0612 0.0714 0.31% 17.29% Average Baseline 0.1003 N/A N/A Sequential BS 0.0965 0.1077 1.30% 4.40%
Sequential BS: Three unknown base stations operating in different time intervals
Local: Post-scale alignment error; Local (Unscaled): Pre-scale alignment error.
Improvement was calculated relative to baseline performance.
Note: V203, AL01, AL03, and AL05 lose tracking before 100 seconds.
Table IV presents results for scenarios where three base stations operate in different time intervals (BS1: 10-40s, BS2: 50-70s, BS3: 80-100s) with unknown locations at frequency 78 GHz as illustrated in Fig. 10. For the monocular configuration, results demonstrate consistent improvement across datasets despite the challenging sequential operation scenario. We observed an average improvement of 4.40% in local accuracy across all sequences, with particularly strong improvements in AL04 (17.29%) and V103 (7.93%). The system maintained reliable scale estimation with an average error of 1.30% across all datasets. These improvements are especially noteworthy considering the lack of continuous base station coverage, suggesting that even intermittent ToA measurements can significantly enhance SLAM performance in minimal sensor configurations.
Figure 10: Bastion Station Activity Over Trajectories: This plot illustrates the active periods of three bastion stations (BS1: 10–40s, BS2: 50–70s, BS3: 80–100s) over the trajectories of various datasets (V101, V102, V103, V201, V202, Dataset 1, and Dataset 3). Each horizontal bar represents the total duration of a dataset, with colored segments highlighting the intervals during which each bastion station is active
V-B2 ToA Integration as an Alternative to Loop Closure
In typical SLAM implementations, loop closure serves as a critical mechanism for drift correction by recognizing previously visited locations. However, this capability faces limitations in many real-world scenarios. Our investigation focuses on whether ToA measurements can provide a viable alternative when traditional loop closure is unavailable or unreliable. This study is conducted using RGBD for the Aerolab dataset and Stereo for the Euroc Mav dataset.
Table V presents a comprehensive comparison between baseline SLAM with loop closure enabled, the same system with loop closure disabled, and two ToA integration configurations using unknown base station locations: sequential operation of three base stations and continuous operation of three base stations. The results reveal a compelling case for ToA integration as a loop closure alternative.
TABLE V: ToA Integration as an Alternative to Loop Closure
Dataset Baseline (LC) Baseline (No LC) ToA (Seq) ToA (Cont) Improvement (Seq / Cont) (%) AL01 0.064 0.153 0.138 0.129 9.8% / 15.7% V102 0.099 0.272 0.205 0.142 24.6% / 47.8% V103 0.161 0.507 0.299 0.257 41.0% / 49.3% V202 0.190 0.213 0.230 0.208 -8.0% / 2.3% V203 0.960 0.832 0.621 0.557 25.4% / 33.1% Average 0.294 0.395 0.299 0.259 18.6% / 29.6%
Baseline (LC): Baseline results with loop closure enabled.
Baseline (No LC): Baseline results with loop closure manually deactivated.
ToA (Seq): Results with 3 sequential base stations (freqseq).
ToA (Cont): Results with 3 continuous base stations (freq78).
Improvement (%) is calculated relative to the baseline without loop closure.
V203 shows higher error due to tracking losses and relocalization.
When loop closure is disabled, we observe significant degradation in localization accuracy across most sequences, with the average ATE increasing from 0.294 m to 0.395 m (a 34.35% degradation). ToA integration effectively mitigates this performance loss, with three sequential base stations improving accuracy by 18.6% compared to the no-loop-closure baseline. The improvement becomes even more pronounced with continuous ToA measurements from three base stations, achieving a 29.6% enhancement and reducing the average error from 0.395 m to 0.259 m.
The improvements are particularly striking in challenging sequences where loop closure is typically critical for maintaining baseline performance. For example, in sequence V104, continuous ToA measurements deliver a remarkable 49.3% improvement in accuracy over the no-loop-closure baseline, while sequence V203 shows a 33.1% improvement. These significant gains demonstrate that ToA integration can effectively replace loop closure in preserving trajectory consistency, even without prior knowledge of base station locations. Interestingly, in sequence V203, the baseline without loop closure achieves a slightly better performance (0.832 meters) compared to the baseline with loop closure (0.960 meters). This counterintuitive result may be attributed to suboptimal pose graph optimizations, particularly in this challenging sequence, which involves tracking losses and relocalization events. A similar pattern is observed in dataset V202, where the baseline with loop closure (0.190 meters) marginally outperforms the baseline without loop closure (0.213 meters). In this case, using ToA as a substitute for loop closure provides only minimal improvements or even degradation.
V-B3 Comparison with UWB-VO and Other State-of-the-Art Methods
TABLE VI: Comparison with UWB-VO and State-of-the-Art Methods on EuRoC Machine Hall Dataset
Dataset UWB-VO VINS-mono OKVINS ROVIO Ours (1BS) Ours (2 Seq BS) Local Unscaled Local Unscaled MH01 0.120 0.270 0.160 0.210 0.102 0.116 0.103 0.129 (1.26%) (1.84%) MH02 0.130 0.150 0.220 0.250 0.119 0.120 0.120 0.125 (0.26%) (0.66%) MH03 0.280 0.130 0.240 0.250 0.208 0.215 0.208 0.212 (1.20%) (0.31%) MH04 0.360 0.250 0.340 0.490 0.289 0.349 0.286 0.317 (2.53%) (1.35%) MH05 0.300 0.350 0.470 0.520 0.300 0.460 0.294 0.375 (3.93%) (1.82%) Mean 0.238 0.230 0.286 0.344 0.204 0.252 0.202 0.232 (1.84%) (1.20%)
1BS: Single base station at [10,10,10] with 5Hz measurements
2BS: Two sequential base stations active during 20-55s and 60-100s respectively
Local: Results with scale correction applied
Unscaled: Results without scale correction
Bold values indicate the best RMSE performance for each sequence
Values in italics show scale errors for our methods
UWB-VO reported scaling errors ¡ 1% but did not provide detailed results
Table VI presents a detailed comparison, introducing both Local (scaled) and Unscaled RMSE values. The Local results reflect the accuracy with scale correction applied (Sim(3) alignment), while the Unscaled results show raw SE(3) alignment-based errors, making direct scale comparisons possible.
The UWB-VO approach, proposed by Li et al., is a tightly coupled SLAM framework that integrates monocular visual odometry with UWB ToA constraints to address scale ambiguity and reduce drift. The system assumes the presence of a single static UWB anchor and was evaluated using the Machine Hall dataset from the EuRoC MAV benchmark. Notably, their experiments rely on simulated UWB ToA data, generated with a Gaussian distribution at a standard deviation of 5 cm and a frequency of 30Hz.
Unlike UWB-VO, which relies on a single simulated UWB anchor at 30 Hz, our approach evaluates both single and sequential base station setups (Fig 11) at a lower, more practical 5 Hz measurement rate. The sequential two-base-station approach improves performance in challenging sequences (MH04, MH05).
Figure 11: Bastion Station Activity Over Trajectories (Euroc MAV Machine Hall): This plot illustrates the active periods of two bastion stations (BS1: 20-55s, BS2: 60-100s) over the trajectories of Machine Hall (MH01-MH05). Each horizontal bar represents the total duration of a dataset, with colored segments highlighting the intervals during which each bastion station is active
Despite using a lower ToA measurement frequency, our method achieves a comparable 0.202 m mean RMSE. Even in Unscaled form, the RMSE remains competitive, demonstrating strong inherent scale estimation without external correction. In contrast, UWB-VO reports ¡1% scale error but does not provide per-sequence details, making direct comparisons difficult.
These results highlight that our sequential base station method and lower measurement rate still yield robust performance, outperforming several established methods while operating under more relaxed and realistic constraints.
VI Analysis of Base Station Configurations for SLAM Performance
In this section, we analyze the impact of the base station position on the SLAM performance. Geometric Dilution of Precision (GDOP) plays a crucial role in determining localization accuracy through optimal base station positioning. In this section, we analyze the impact of base station configuration on localization performance. GDOP quantifies how the geometric arrangement of base stations affects position estimation accuracy and is calculated as:
$G=\begin{bmatrix}\frac{x_{1}-x_{r}}{d_{1}}&\frac{y_{1}-y_{r}}{d_{1}}&\frac{z_{1}-z_{r}}{d_{1}}&1\\ \frac{x_{2}-x_{r}}{d_{2}}&\frac{y_{2}-y_{r}}{d_{2}}&\frac{z_{2}-z_{r}}{d_{2}}&1\\ \vdots&\vdots&\vdots&\vdots\\ \frac{x_{n}-x_{r}}{d_{n}}&\frac{y_{n}-y_{r}}{d_{n}}&\frac{z_{n}-z_{r}}{d_{n}}&1\end{bmatrix}$ (7)
$\text{GDOP}=\sqrt{\text{trace}((G^{T}G)^{-1})}$ (8)
Lower GDOP values, achieved when a mobile device is surrounded by well-distributed base stations, correspond to higher positioning accuracy, while clustered configurations result in higher GDOP and degraded performance. Our analysis examines five representative base station arrangements (Tetrahedral, Asymmetric, Diamond, Z-Shape, and Clustered) to demonstrate how geometric configuration directly impacts localization precision, particularly in applications requiring high accuracy such as drone navigation. For a deeper exploration of GDOP and its optimization, readers are referred to [44] and [45].
we use five distinct base station configurations (Fig.12) with exact coordinates detailed in Table VII:
Tetrahedral: 3D distributed (ceiling center + ground triangle)
Z-Shape: Diagonally alternating heights
Asymmetric: Realistic irregular placement
Diamond: Distributed evenly at two different heights
Clustered: Single-wall co-planar arrangement
(a) Tetrahedral
(b) Z-shaped
(c) Asymmetric
(d) Diamond
(e) Clustered
Figure 12: Five configurations of base stations (BS1-BS4) with trajectory data from the EuRoC MAV V101 dataset. Top row: 3D representations; Bottom row: corresponding 2D X-Y projections with z-coordinates. Configurations are ordered by performance (best to worst): Tetrahedral (ATE: 0.111 m), Z-shaped (0.132 m), Asymmetric (0.141 m), Corner (0.161 m), and Clustered (0.182 m).
TABLE VII: Exact 3D coordinates (in meters) of base stations for all evaluated configurations.
Base Station Tetrahedral Diamond Z-Shape Asymmetric Clustered BS1 (0, 0, 3) (0, -5, 0.5) (-5, -5, 0.5) (-5, -2, 3) (-5, -3, 1) BS2 (-4, -4, 0) (5, 0, 1) (5, -5, 3) (2, -5, 1) (-5, -1, 2) BS3 (4, -4, 0) (0, 5, 0.5) (-5, 5, 3) (5, 3, 2) (-5, 1, 1) BS4 (0, 4, 0) (-5, 0, 1) (5, 5, 0.5) (-1, 5, 0.5) (-5, 3, 2)
These configurations were evaluated across both EuRoC MAV and Aerolab Dataset for Stereo and RGB-D respectively.
VI-A Results on EuRoC MAV Dataset
Table VIII presents the average ATE values for both local and global accuracy across all configurations and EuRoC MAV dataset sequences.
TABLE VIII: ATE for different base station configurations across EuRoC MAV datasets (meters). Each cell contains local accuracy / global accuracy values.
Configuration V101 V102 V103 V201 V202 Average Tetrahedral 0.078/0.083 0.124/0.135 0.140/0.158 0.124/0.158 0.184/0.189 0.130/0.145 Z-Shape 0.079/0.082 0.108/0.129 0.177/0.203 0.087/0.135 0.188/0.199 0.128/0.150 Asymmetric 0.079/0.083 0.105/0.118 0.186/0.239 0.131/0.207 0.183/0.203 0.137/0.170 Diamond 0.078/0.083 0.186/0.286 0.148/0.162 0.084/0.118 0.177/0.188 0.135/0.167 Clustered 0.084/0.095 0.137/0.155 0.223/0.628 0.096/0.245 0.189/0.204 0.146/0.265
The results demonstrate the impact of the base station geometry on both local and global SLAM accuracy, with the Z-Shape configuration showing the best average local accuracy (0.128 m) and the Tetrahedral configuration achieving the best average global accuracy (0.145 m). The Clustered configuration showed the poorest performance in both metrics, with a particularly significant degradation in global accuracy (0.265 m, 82.8% worse than the Tetrahedral configuration).
VI-B Results on AeroLab Dataset
The same evaluation of different base station configurations was conducted using the AeroLab dataset. Table IX presents the average ATE values.
TABLE IX: ATE for different base station configurations with 5G ToA integrated RGB-D across AeroLab datasets (meters). Each cell contains local accuracy / global accuracy values.
Configuration AL01 AL02 AL03 AL04 AL05 Average Tetrahedral 0.063/0.092 0.110/0.115 0.143/0.150 0.092/0.095 0.165/0.166 0.115/0.124 Z-Shape 0.062/0.095 0.107/0.125 0.146/0.155 0.095/0.105 0.157/0.165 0.113/0.129 Asymmetric 0.068/0.112 0.109/0.142 0.155/0.160 0.099/0.103 0.172/0.176 0.121/0.138 Diamond 0.062/0.097 0.110/0.138 0.143/0.152 0.097/0.117 0.165/0.172 0.115/0.135 Clustered 0.069/0.131 0.113/0.144 0.148/0.159 0.091/0.096 0.176/0.184 0.119/0.143
The AeroLab results show almost the same pattern as the EuRoC dataset. The Z-Shape configuration achieved the best average local accuracy (0.113 m), closely followed by the Tetrahedral and Diamond configurations (both 0.115 m). For global accuracy, the Tetrahedral configuration maintained its advantage (0.124 m), with the Z-Shape configuration a close second (0.129 m). The Clustered configuration again showed the poorest performance overall, particularly in global accuracy (0.143 m).
VI-C Cross-Dataset Analysis
Comparing the results across both datasets reveals several consistent patterns and some interesting differences, as shown in Table X.
TABLE X: Combined average RMSE across both datasets (meters)
Configuration EuRoC Local/Global AeroLab Local/Global Combined Local/Global Tetrahedral 0.130/0.145 0.115/0.124 0.123/0.135 Z-Shape 0.128/0.150 0.113/0.129 0.121/0.140 Diamond 0.135/0.167 0.115/0.135 0.125/0.151 Asymmetric 0.137/0.170 0.121/0.138 0.129/0.154 Clustered 0.146/0.265 0.119/0.143 0.133/0.204
The combined results highlight several key insights:
  • The Z-Shape configuration shows the best overall local accuracy (0.121m), with the Tetrahedral configuration a close second (0.123m).
  • The Tetrahedral configuration delivers the best overall global accuracy (0.135m), outperforming other configurations by at least 3.7%.
  • The Clustered configuration performs consistently worse across both datasets, with significantly degraded global accuracy (0.204m, 51.1% worse than Tetrahedral).
  • The performance gap between configurations is more pronounced in the EuRoC dataset than in the AeroLab dataset, suggesting that environment complexity and motion dynamics play important roles in determining the impact of base station geometry.
VI-D Global-to-Local Accuracy Ratio
The ratio between global and local accuracy provides insight into how well each configuration maintains global consistency. Table XI presents these ratios for both datasets.
TABLE XI: Global-to-local accuracy ratio for different base station configurations
Configuration EuRoC Ratio AeroLab Ratio Combined Ratio
Tetrahedral 1.12 1.08 1.10
Z-Shape 1.17 1.14 1.16
Diamond 1.24 1.17 1.21
Asymmetric 1.24 1.14 1.19
Clustered 1.82 1.20 1.53
The Tetrahedral configuration shows the lowest ratios across both datasets (1.10 combined), indicating that its local trajectory accuracy translates well to global positioning. The Clustered configuration has the highest ratio (1.53 combined), demonstrating significant degradation when moving from local to global accuracy assessment. This degradation is particularly severe in the EuRoC dataset (ratio of 1.82), highlighting the vulnerability of the Clustered configuration to challenging conditions.
VI-E Relationship Between GDOP and SLAM Performance
The bar chart in Figure 13 and 14 presents the average GDOP values for each base station configuration across all EuRoC MAV dataset sequences and Aerolab respectively, providing direct empirical evidence of how geometric arrangement affects positioning quality. A correlation between GDOP values and SLAM performance across both datasets is evident especially for global accuracy. The Tetrahedral configuration’s superior global accuracy directly corresponds to its lowest average GDOP in both datasets. However, the fusion with visual data helps mitigate some of the geometric limitations, especially for local trajectory accuracy, resulting in smaller percentage differences between configurations compared to pure localization.
Figure 13: Comparison of average Geometric Dilution of Precision (GDOP) values across different base station configurations for all EuRoC MAV dataset sequences. Lower GDOP values indicate better geometric configuration for localization.
Figure 14: Comparison of average Geometric Dilution of Precision (GDOP) values across different base station configurations for all AeroLab dataset sequences. Lower GDOP values indicate better geometric configuration for localization.
VII System Limitations and Future Enhancements
A fundamental consideration of this work, though inherent to ToA-based positioning rather than our framework’s design, is the assumption of fixed and precisely known base station positions. Firstly, this assumption is crucial for enabling global SLAM capabilities, as these stations serve as critical global references. Global SLAM fundamentally relies on at least some fixed and known reference points to establish absolute positioning. While it is theoretically possible to estimate base station positions dynamically, such approaches typically confine SLAM to local frames, thereby limiting its global applicability. Even if we consider only local SLAM and neglect global positioning requirements, another challenge emerges: Unlike traditional SLAM landmarks, which provide rich 3D or SE3 (6-DoF) measurements and can self-correct during optimization, ToA measurements are scalar and offer only radial distance constraints. This fundamental difference creates observability challenges, as each ToA reading constrains the base station position to a sphere around the UAV. While multiple measurements along a trajectory create intersecting spheres that could theoretically pinpoint base station positions, in practice ToA measurement accuracy (typically 20-30 cm) expands each sphere into a thick shell. This measurement uncertainty, rather than geometric ambiguity, is what makes reliable convergence challenging without accurate initial position estimates for both the drone’s pose and base station positions. This significantly complicates system initialization and contrasts with the current framework, which assumes that the drone’s initial pose is unknown.
In practical deployment scenarios, this limitation is mitigated by the nature of indoor 5G infrastructure, which typically involves fixed installations and one-time deployment. Calibrating base station positions during setup ensures consistent global localization while significantly reducing system complexity. While some applications may involve dynamic or unknown base station positions, maintaining a subset of known positions as global anchors remains essential for establishing a consistent global reference frame. This trade-off between practicality and flexibility enables robust global SLAM while addressing key challenges associated with scalar ToA measurements.
Beyond the base station positioning challenges, several other limitations affect the system’s real-world applicability. The current ToA simulation framework assumes ideal Line-of-Sight (LoS) conditions, while Non-Line-of-Sight (NLoS) conditions, caused by obstacles and reflections, can significantly degrade ToA measurement accuracy and impact SLAM performance. This limitation is compounded by the use of static information matrices for both ToA and visual-inertial measurements, which fail to capture the dynamic nature of measurement uncertainties in real environments. Moreover, while simulation provides a controlled testing environment, it may not fully capture the complexities of real-world signal propagation, hardware noise, and environmental dynamics. These limitations highlight the need for future validation with experimental ToA measurements and the development of more sophisticated uncertainty models that can adapt to changing environmental conditions.
VIII Conclusion
This paper presented a novel approach for integrating 5G Time of Arrival (ToA) measurements into ORB-SLAM3, enabling globally consistent localization and mapping for indoor drone navigation. By extending ORB-SLAM3’s optimization pipeline to jointly process ToA data alongside visual and inertial measurements, we successfully transformed inherently local SLAM estimates into globally referenced trajectories while effectively resolving scale ambiguity in monocular configurations. Our comprehensive evaluation across multiple datasets and SLAM configurations demonstrated that ToA integration maintains high local accuracy while enabling consistent global positioning - a capability absent in traditional SLAM systems. For monocular systems specifically, the integration of ToA measurements successfully addressed the fundamental challenge of scale ambiguity without requiring additional sensors or manual calibration.
The investigation into unknown base station positions revealed that our approach remains effective even with limited knowledge of anchor positions, showcasing the system’s adaptability to real-world deployment constraints. Furthermore, we demonstrated that ToA measurements can serve as a viable alternative to traditional loop closure mechanisms, maintaining trajectory consistency in scenarios where loop detection might be unreliable or impossible. Our analysis of different geometric arrangements of base stations provided valuable insights into optimal deployment strategies, revealing performance differences between configurations. This knowledge directly informs practical implementation by guiding base station placement decisions.
Comparative analysis with state-of-the-art methods confirmed the robustness of our approach even under challenging operational constraints such as lower measurement frequencies and sequential base station operation. The experimental results collectively validate that 5G ToA integration offers substantial benefits for global SLAM applications, particularly in challenging indoor environments where accurate positioning is critical. Future work could explore real-world implementation with physical 5G infrastructure, investigate dynamic base station calibration techniques, and extend the approach to multi-agent collaborative mapping scenarios.